# DeepCompile: A Compiler-Driven Approach to Optimizing Distributed Deep Learning Training

### v1

@@ Abstract.
The increasing scale of deep learning models has led to the development of various parallelization strategies for distributed training across accelerators. For example, fully sharded approaches like DeepSpeed ZeRO-3 and FSDP partition the parameters of each layer across multiple GPUs and gather them through communication when needed. These methods rely on optimizations such as prefetching, which initiates communication early to overlap it with computation and reduce communication overhead, and unsharding, which retains as many parameters in their unsharded form as possible to reduce communication volume. Although the timing of prefetching should be adjusted in response to dynamic memory usage during execution, these systems lack the flexibility to control it, which limits the benefits of prefetching. Moreover, they cannot anticipate how memory usage will change after prefetching is applied, making it difficult to combine it effectively with other optimizations such as unsharding. We present DeepCompile, which compiles user-defined models into computation graphs and applies a sequence of profiling-guided optimization passes for distributed training. Taking dynamic memory usage into account, these passes flexibly insert, reorder, or remove operations to improve communication–computation overlap, reduce memory pressure, and coordinate multiple optimizations in a unified manner. To evaluate the effectiveness of this design, we implemented a fully sharded approach like ZeRO-3 and FSDP on top of DeepCompile, along with three optimizations: proactive prefetching, selective unsharding, and adaptive offloading. We evaluate DeepCompile on the training of Llama 3 70B and Mixtral 8×7B MoE models. DeepCompile achieves up to 1.28× and 1.54× performance improvements over ZeRO-3 and FSDP baselines, respectively, and up to a 7.01× throughput increase in settings with limited GPU resources, using offloading.
@@ 1. Introduction
The rapid growth of deep learning has led to the emergence of increasingly large models. Modern architectures often contain billions of parameters, resulting in immense computational demands. In this context, efficient parallelization across multiple accelerator devices, such as GPUs, has become essential for achieving feasible training times and manageable costs.
@@
There are several parallelization strategies for training large models: In data parallelism, all model parameters are replicated across multiple GPUs. This approach is natively supported in many deep learning frameworks, including PyTorch (Paszke et al., 2019). Pipeline parallelism (Huang et al., 2019) instead splits the model by layers, assigning each block to a different GPU. Tensor parallelism (Shoeybi et al., 2019) divides parameter tensors within a layer across GPUs, where each device computes a partial result that is later aggregated. In addition to these, the fully sharded approach, which is implemented in DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) (hereafter ZeRO-3) and Fully Sharded Data Parallel (FSDP) (Team, 2023), partitions the parameters of each layer across multiple GPUs and gathers them through communication before each layer’s computation begins.
While these strategies provide scalability, efficiently managing memory and overlapping communication with computation to reduce communication overhead (Wang et al., 2022, 2024) remain open challenges. Existing systems rely on various optimization techniques to address these issues. In the context of the fully sharded approach, such as ZeRO-3 and FSDP, prefetching initiates communication early to gather partitioned parameters and overlap it with computation; unsharding keeps parameters in their full form to reduce communication when memory permits; and offloading, which moves data from GPU memory to host memory, is often used in combination when GPU memory is insufficient.
Although these optimizations are effective in principle, applying them efficiently in practice is not straightforward. Take prefetching as an example. In the fully sharded approach, parameters partitioned across multiple GPUs must be gathered via all-gather communication, which requires a large buffer to hold the gathered parameters. Initiating the all-gather earlier increases opportunities to overlap communication with computation. However, it also extends the buffer’s lifetime and increases memory pressure, since the buffer cannot be released until all computations that use the parameters have completed. Similarly, the overhead of offloading can be mitigated by minimizing data transfers between host and GPU memory and overlapping them with computation. To make these optimizations effective, it is essential to flexibly adjust the timing of all-gather and offloading operations, monitor memory usage trends, and analyze data dependencies across broad scopes such as the forward and backward passes.
Moreover, when multiple optimizations are applied together, their interactions must be carefully coordinated. For instance, unsharding can improve performance when communication is on the critical path, but when combined with prefetching, it must also account for the additional memory pressure introduced by early optimization.
In fully sharded approaches like ZeRO-3 and FSDP, all-gather operations are implemented via runtime hooks inserted into the user-level code (typically in Python). Optimizations such as prefetching and unsharding are supported, but this design lacks the flexibility to adjust the timing of all-gather or offloading operations based on runtime memory usage over broader scopes, such as an entire forward or backward pass. Moreover, there is no clear mechanism to coordinate multiple optimizations in a unified manner.
Other systems have been proposed to automatically combine multiple optimizations for distributed training based on profiling. Examples include Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), and Unity (Unger et al., 2022), which aim to determine effective combinations of parallelization strategies such as data, tensor, and pipeline parallelism, as well as to apply kernel-level optimizations like kernel fusion. However, these systems primarily focus on static planning of parallelization strategies and kernel-level optimizations. They do not take into account runtime dynamics such as changes in memory usage during execution. As a result, they lack mechanisms to adjust communication schedules, adapt to memory availability at runtime, or coordinate optimizations such as prefetching, unsharding, and offloading.
To address these limitations, we introduce DeepCompile, a compiler-based system that uses an existing compiler to convert the model into a computation graph and then transforms it for optimization. DeepCompile applies a series of optimization passes, where each pass targets a specific goal—such as improving communication overlap—by inserting, removing, or reordering operations in the graph. These passes make decisions based on profiling such as operator execution time and memory usage trends over broader scopes, including entire forward and backward passes. Since each pass produces a modified graph, subsequent passes can analyze the result and flexibly adapt to the effects of earlier transformations.
@@
To evaluate the effectiveness of this design, we implemented the fully sharded approach, which is similar to ZeRO-3 and FSDP, as a series of optimization passes within DeepCompile. In addition, we implemented the following key optimizations as separate passes.
Proactive prefetching.:
To maximize overlap between communication and computation, this optimization pass initiates all-gather as early as possible, considering how available memory changes as the forward and backward passes progresses.
Selective unsharding.:
This pass keeps as many parameters unsharded as possible to reduce communication overhead caused by all-gather communication, and decides which parameters to unshard based on operator-level memory profiling.
Adaptive offloading.:
@@
DeepCompile offloads optimizer states such as momentum and variance used by Adam (Kingma and Ba, 2014) to CPU memory when GPU memory is insufficient. To reduce data transfer overhead, it offloads only the amount of data that exceeds the memory limit and schedules transfers to overlap with computation.
SimpleFSDP (Zhang et al., 2024) takes a similar approach to DeepCompile by using compilation to optimize the placement of all-gather operations in the fully sharded approach. However, it focuses specifically on prefetching and does not address other optimizations such as unsharding or offloading, which leads to notable differences in design. In SimpleFSDP, communication operations are inserted into the user-level code (Python) before compilation, whereas DeepCompile first compiles the model into a computation graph and then applies transformations. In SimpleFSDP’s design, the compiler must correctly lower these manually inserted communication operations into graph-level operators. While the PyTorch compiler currently supports lowering all-gather operations used by SimpleFSDP, it may fail to handle other optimizations, such as offloading, if they are implemented at the user level. This design limits the extensibility of SimpleFSDP in supporting and combining a broader range of optimizations.
@@
We evaluated DeepCompile on large-scale distributed training tasks using Llama 3 70B and Mixtral 8x7B MoE models. Across all configurations, DeepCompile consistently improved training efficiency compared to baselines, including ZeRO-3, FSDP, and setups using PyTorch compiler optimizations. It achieved up to 1.28× improvement on Llama 3 70B and up to 1.54× on the Mixtral 8x7B MoE model. Additionally, it demonstrated a 7.01× increase in throughput when using fewer GPUs with offloading.
@@ 2. Background and Motivating Example
2.1. Distributed Training
The emergence of Transformer architectures (Vaswani et al., 2017), which enabled breakthroughs like BERT (Devlin et al., 2019), has significantly accelerated the growth of modern deep learning models, such as GPT-3 (Brown et al., 2020), Llama-2 (Touvron et al., 2023), Llama-3 (Grattafiori et al., 2024), Phi-3 (Abdin et al., 2024), PaLM (Chowdhery et al., 2022), and OPT (Zhang et al., 2022). This trend has led to a tremendous increase in computational demands. While accelerator devices such as GPUs have become the standard, cutting-edge models now require distributed training across multiple GPUs to be feasible. For example, Llama-3 (Grattafiori et al., 2024) was trained on 16,384 NVIDIA H100 GPUs over 54 days. Even fine-tuning, which adapts pre-trained models to specific downstream tasks, often involves dozens of GPUs running for hours or days.
In distributed training, where multiple GPUs are used in parallel, reducing overhead from both inter-device communication and memory management is critical for efficiently training large-scale models.
@@ 2.2. Parallelization Strategies
Various parallelization strategies have been proposed to improve the efficiency of distributed training.
Data parallelism replicates the entire model on each GPU, and each GPU processes a different portion of the training data. A key advantage is that it can be used with any model architecture without requiring code changes, and the communication overhead is relatively low. However, because each GPU must store a full copy of the model, large models may exceed the memory capacity.
Unlike data parallelism, pipeline parallelism partitions the model to fit within GPU memory by grouping layers into coarser-grained stages and placing each stage on a different GPU. The training batch is split into smaller fragments and processed in a pipelined fashion. To fully utilize all GPUs, the compute time of each stage should be approximately balanced. In addition, the choice of partition points affects the communication volume between GPUs. These factors often require manual tuning to achieve optimal performance.
Tensor parallelism is primarily applied to the multi-head attention and MLP layers in Transformer models. It splits specific parameter tensors along designated dimensions across GPUs, performs computations locally, and aggregates results using all-reduce communication. For example, in an MLP block, the first weight matrix is typically split horizontally, and the second one vertically; each GPU computes partial results, and an all-reduce is applied after the second linear transformation to combine them. In addition, although well-tested implementations exist for standard architectures such as GPT (Brown et al., 2020), adapting tensor parallelism to model variants such as multi-query attention (MQA) (Shazeer, 2019), grouped-query attention (GQA) (Narang et al., 2021), or multi-head latent attention (Liu et al., 2024) requires additional engineering effort. In each case, developers must manually identify which parameters can be split, choose appropriate dimensions for partitioning, and ensure correct synchronization across devices.
An alternative to pipeline and tensor parallelism is the fully sharded approach, which partitions the parameters of each layer across multiple GPUs. Frameworks such as Fully Sharded Data Parallel (FSDP) and ZeRO-3 adopt this strategy. During training, each GPU gathers its required parameter shards using all-gather communication before computing each layer. To reduce peak memory usage, the gathered parameters are discarded immediately after the layer’s computation finishes. This approach is widely adopted because it can be applied to arbitrary model architectures without requiring manual restructuring. In this work, we focus on optimizing this fully sharded approach.
@@ 2.3. Deep Learning Compiler
Deep learning compilers have rapidly advanced in recent years (Chen et al., 2018; Google, 2023; Ansel et al., 2024). They translate high-level models into computation graphs, which enables graph-level analysis and flexible optimizations such as operator reordering, fusion, and memory reuse.
@@ 2.4. Motivating Example
The fully sharded approach, as implemented in systems like ZeRO-3 and FSDP, employs runtime optimizations such as prefetching and unsharding. However, these optimizations alone are not sufficient to fully improve training efficiency. Let us take prefetching as an example. Prefetching aims to reduce communication overhead by initiating all-gather operations earlier than the layer where the parameters are actually needed, thereby overlapping communication with computation. Because the fully sharded approach allocates a large buffer to store the gathered parameters, initiating all-gather earlier extends the buffer’s lifetime and increases memory pressure. Therefore, prefetching needs to be scheduled based on memory usage patterns throughout the forward and backward passes.
@@
Figure 1 illustrates a typical pattern of GPU memory usage during a forward and backward pass. Memory usage gradually increases during the forward pass as activations are stored, and decreases during the backward pass as activations are released. At the beginning of the forward pass, a substantial amount of unused memory is available, providing an opportunity to prefetch aggressively and increase computation-communication overlap. In contrast, near the end of the forward pass and throughout the backward pass, memory usage is high, limiting the opportunity for prefetching due to reduced memory availability. As the backward pass progresses and memory is gradually freed, more aggressive prefetching becomes possible again. Existing systems such as ZeRO-3 and FSDP typically allow users to specify a fixed buffer size for prefetching. However, this static configuration cannot adapt to changes in available memory as computation progresses, limiting the potential benefits of prefetching.
@@
Figure 1. Memory usage trends and scheduling opportunities for prefetching and offloading. Profile of several final layers with a sequence length of 4096 and a vocabulary size of 128k. Significant memory spikes are observed in the log-softmax and negative log-likelihood loss layers.
@@
The same issue arises with offloading, which transfers data such as optimizer states (e.g., momentum and variance in Adam optimizer (Kingma and Ba, 2014)) to host memory to reduce GPU memory usage. While offloading can significantly reduce peak GPU memory, the associated data transfers introduce overhead. A naive strategy is to offload states before starting the forward pass, but this leads to significant overhead due to idle GPU time. Instead, data transfers and computation can be overlapped. By initiating transfers at the beginning of the forward pass and synchronizing their completion just before memory usage rises to its limit, it is possible to hide the cost of data movement. In the backward pass, the opposite strategy can be applied: as memory usage decreases, transfers from host memory to GPU can be initiated.
@@
These observations highlight the limitations of existing approaches. To address these limitations, we introduce DeepCompile, a compiler-based approach designed to systematically optimize fully sharded training.

### v2

@@ Abstract.
The rapid growth of deep learning models has increased the demand for efficient distributed training strategies. Fully sharded approaches like ZeRO-3 and FSDP partition model parameters across GPUs and apply optimizations such as prefetching and unsharding to reduce communication overhead. However, these systems lack fine-grained control over memory and communication scheduling, making it difficult to balance computation–communication overlap with memory requirements. Coordinating multiple optimizations such as prefetching and unsharding is also difficult, since their effects on memory usage can influence each other. To tackle these challenges, we propose DeepCompile, a compiler-based optimization framework for distributed training. DeepCompile transforms user-defined models into computation graphs and applies a series of profiling-guided optimization passes, each modifying the graph based on profiling information such as execution time and memory usage. This design allows each pass to flexibly insert, reorder, or remove operations such as all-gather and memory allocation, improving communication–computation overlap and reducing memory pressure. Each pass can access updated profiling feedback from earlier passes, enabling coordinated optimizations. We further enhance DeepCompile by three additional optimizations: proactive prefetching, selective unsharding, and adaptive offloading. Our evaluation shows that DeepCompile achieves up to 1.28× and 1.54× speedups over ZeRO-3 and FSDP baselines, respectively, and up to a 7.01× throughput increase in settings with limited GPU resources using offloading.
@@ 1. Introduction
The rapid growth of deep learning has led to the emergence of increasingly large models. Modern architectures often contain tens or even hundreds of billions of parameters (Brown et al., 2020; Zhang et al., 2022; Abdin et al., 2024; Grattafiori et al., 2024; Chowdhery et al., 2022), resulting in immense computational demands. While deep learning compilers (Chen et al., 2018; Google, 2023; Ansel et al., 2024) have been highly successful in improving training efficiency on a single accelerator device, such as a GPU, efficient parallelization across multiple devices has become essential to fit large models into memory.
@@
There are several parallelization strategies for training large models. In data parallelism, all model parameters are replicated across multiple GPUs. This approach is natively supported in many deep learning frameworks, including PyTorch (Paszke et al., 2019). Pipeline parallelism (Huang et al., 2019) instead splits the model by layers, assigning each block to a different GPU. Tensor parallelism (Shoeybi et al., 2019) divides parameter tensors within a layer across GPUs, where each device computes a partial result that is later aggregated. In addition to these, the fully sharded approach, which is implemented in DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) (hereafter ZeRO-3) and Fully Sharded Data Parallel (FSDP) (Team, 2023), partitions the parameters of each layer across multiple GPUs and gathers them through all-gather communication before each layer’s computation begins.
Although existing parallelization strategies have enabled large-scale model training, two key challenges remain unsolved:
Challenge 1: Fine-grained control over communication and memory scheduling. Existing systems often lack the ability to precisely control when to initiate communication or allocate and release memory buffers. As a result, it is difficult to optimize the overlap between communication and computation, or to efficiently manage memory usage that fluctuates during execution.
Challenge 2: Coordinating multiple optimizations. As existing deep learning compilers apply a sequence of independent optimization passes, such as kernel fusions and operator reordering, distributed training also requires independent optimizations such as communication scheduling and memory planning. However, there is currently no unified mechanism to combine these optimizations and handle their interactions effectively.
To illustrate Challenge 1, consider prefetching in the fully sharded approach. Initiating all-gather communication earlier can improve overlap with computation, but prefetching too much data can exhaust available memory, as the buffers cannot be released until all dependent computations are completed.
A representative case of Challenge 2 involves combining prefetching with unsharding, which retains as many parameters in their unsharded form as possible to reduce communication overhead. To combine these two optimizations effectively, the system must accurately estimate how much memory remains available after prefetching is applied.
Existing systems partially address these challenges. Runtime-based approaches such as ZeRO-3 and FSDP support optimizations like prefetching and unsharding, but lack fine-grained control over the timing of communication and memory scheduling in response to runtime memory changes. Compiler-based systems such as Alpa (Mo et al., 2022), Unity (Unger et al., 2022), nnScalar (Lin et al., 2024) automatically combine different parallelization strategies, but do not adapt to dynamic memory behavior during execution.
To bridge this gap, we introduce DeepCompile, a compiler-based optimization framework for distributed training. DeepCompile uses an existing compiler to convert a model into a computation graph and applies a series of transformations to optimize multi-GPU execution. These transformations improve communication overlap and reduce memory pressure based on profiling information that captures execution time and memory usage trends across the forward and backward passes.
This approach addresses the two key challenges discussed earlier:
Graph transformations enabling fine-grained memory-aware scheduling. By directly transforming the computation graph, DeepCompile can flexibly inject, reorder, and remove operations such as all-gather. This fine-grained control enables precise placement of communication and memory operations, which is difficult to express in user-level code. Furthermore, by capturing operator-level memory usage patterns across forward and backward passes, DeepCompile identifies optimal points to initiate communication and allocate or release memory based on runtime memory availability.
Composable and unified optimization passes with profiling feedback. DeepCompile organizes optimizations into modular optimization passes, each transforming the computation graph based on profiling data such as execution time and memory usage. While each pass can operate sequentially, later passes can access updated profiling data from the transformed graph to observe and respond to the effects of earlier passes. This design enables multiple distributed training optimizations to be composed in a globally coordinated manner.
@@
To evaluate the effectiveness of this design, we implemented the fully sharded approach in DeepCompile as an optimization pass, similar in functionality to ZeRO-3 (Rajbhandari et al., 2020) and FSDP (Team, 2023)4.1). This pass demonstrates that the fully sharded paradigm can be captured within the optimization pass framework, and serves as a foundation for building additional compiler-based optimizations.
Building on this foundation, we design and incorporate three additional optimization passes into DeepCompile. Each pass targets a distinct aspect of communication or memory optimization in fully sharded training.
  • Proactive prefetching (§4.2) initiates all-gather operations earlier to improve overlap between communication and computation, while taking into account dynamic memory availability during execution.
  • Selective unsharding (§4.3) avoids sharding for selected parameters based on operator-level memory profiling, reducing all-gather operations for sharded parameters and adapting to memory usage trends.
@@
  • Adaptive offloading (§4.4) transfers optimizer states, such as the momentum and variance buffers used by the Adam optimizer (Kingma and Ba, 2014), to CPU memory when GPU memory is insufficient. These transfers are scheduled to overlap with computation to hide latency.
All of these passes are applied after the optimization pass enabling the baseline fully sharded functionality described in §4.1. Moreover, as we show in later experiments, combining proactive prefetching and selective unsharding yields significant performance improvements over applying either pass in isolation.
SimpleFSDP (Zhang et al., 2024) takes a similar approach to DeepCompile by using compilation to improve the placement of all-gather operations in the fully sharded approach. However, it is specialized for FSDP and focuses solely on prefetching, along with fused all-gather communications. As a result, it does not support composing multiple modular optimization passes. Coordination across passes using profiling feedback—for example, adapting later optimizations to the effects of earlier ones—is not considered in its design.
@@
We evaluated DeepCompile on large-scale distributed training tasks using Llama 3 70B and Mixtral 8x7B MoE models. Across all configurations, DeepCompile consistently improved training efficiency compared to baselines, including ZeRO-3, FSDP, and setups using PyTorch compiler optimizations. It achieved up to 1.28× improvement on Llama 3 70B and up to 1.54× on the Mixtral 8x7B MoE model. Additionally, it delivered a 7.01× increase in throughput in low-GPU settings where, even with parameter partitioning, the model does not fit in GPU memory without offloading.
@@ 2. Background and Motivating Example
The rapid scaling of Transformer-based architectures (Vaswani et al., 2017) has driven a dramatic increase in model size and computational demands. While accelerators like GPUs are standard, training state-of-the-art models now requires large-scale distributed training. For instance, Llama-3 (Grattafiori et al., 2024) was trained on 16,384 NVIDIA H100 GPUs over 54 days. Even fine-tuning tasks frequently span dozens of GPUs. As such, reducing overhead from inter-device communication and memory usage has become essential for efficient training.
@@ 2.1. Parallelization Strategies
To address the computational and memory demands of large-scale models, several parallelization strategies have been developed.
Data parallelism replicates the entire model on each GPU and splits the input data. While easy to adopt, it cannot scale to very large models that exceed single-GPU memory. Pipeline parallelism splits the model by layers into sequential stages across GPUs. It reduces memory usage but requires careful balancing of stage workloads and often manual tuning. Tensor parallelism partitions individual weight matrices across devices and aggregates partial results using all-reduce. Though effective for standard architectures with well-supported implementations (e.g., GPT (Brown et al., 2020)), it often requires non-trivial engineering to adapt to variants such as multi-query attention (Shazeer, 2019) or grouped-query attention (Ainslie et al., 2023).
An alternative to pipeline and tensor parallelism is the fully sharded approach, which partitions the parameters of each layer across multiple GPUs. Frameworks such as DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) and Fully Sharded Data Parallel (FSDP) (Team, 2023) adopt this strategy. In this approach, each GPU gathers its required parameter shards using all-gather communication before computing each layer. To reduce peak memory usage, the gathered parameters are discarded immediately after the layer’s computation finishes. This approach is widely adopted because it can be applied to arbitrary model architectures without requiring manual restructuring. In this work, we focus on optimizing this fully sharded approach.
@@ 2.2. Deep Learning Compiler
Deep learning compilers have rapidly advanced in recent years (Chen et al., 2018; Google, 2023; Ansel et al., 2024). They translate user-defined models into computation graphs, which enables graph-level analysis and flexible optimizations such as operator reordering, fusion, and memory reuse.
@@ 2.3. Motivating Example
The fully sharded approach, as implemented in systems like ZeRO-3 and FSDP, employs runtime optimizations such as prefetching and unsharding. However, these optimizations alone are not sufficient to fully improve training efficiency.
@@
Let us take prefetching as an example. Prefetching aims to reduce communication overhead by initiating all-gather operations earlier than the layer where the parameters are actually needed, thereby overlapping communication with computation. Because the fully sharded approach allocates a large buffer to store the gathered parameters, initiating all-gather earlier extends the buffer’s lifetime and increases memory pressure. Therefore, prefetching needs to be scheduled based on memory usage patterns throughout the forward and backward passes.
@@
Figure 1 illustrates a typical pattern of GPU memory usage during forward and backward passes. Memory usage gradually increases during the forward pass as activations are stored for the following backward pass, and decreases during the backward pass as activations are released. At the beginning of the forward pass, a substantial amount of unused memory is available, providing an opportunity to prefetch aggressively and increase computation-communication overlap. In contrast, near the end of the forward pass and throughout the backward pass, memory usage is high, limiting the opportunity for prefetching due to reduced memory availability. As the backward pass progresses and memory is gradually freed, more aggressive prefetching becomes possible again. Existing systems such as ZeRO-3 and FSDP typically allow users to specify a fixed buffer size for prefetching. However, this static configuration cannot adapt to changes in available memory as computation progresses, limiting the potential benefits of prefetching.
@@
Figure 1. Memory usage trends and scheduling opportunities for prefetching and offloading. Profile of several final layers with a sequence length of 4096 and a vocabulary size of 128k. Significant memory spikes are observed in the log-softmax and negative log-likelihood loss layers in the backward pass.
@@
The same issue arises with offloading, which transfers data such as optimizer states (e.g., momentum and variance in the Adam optimizer (Kingma and Ba, 2014)) to host memory to reduce GPU memory usage during forward and backward passes. This is possible because optimizer states are only needed after the backward pass, when updating parameters. Though offloading can significantly reduce peak GPU memory usage, the associated data transfers introduce overhead. A naive strategy is to offload optimizer states before the forward pass and reload them after the backward pass finishes, since they are not used during computation. However, this leads to idle GPU time during data transfers.
Instead, data transfers and computation can be overlapped. By initiating transfers at the beginning of the forward pass and synchronizing their completion just before memory usage reaches its peak, it is possible to hide the cost of data movement. In the backward pass, the opposite strategy can be applied: as memory usage decreases, transfers from host memory to GPU can be initiated. To achieve this, the system must accurately track how memory usage changes during execution and place offload, reload, and synchronization operations at appropriate points.
@@
These observations motivate the design of DeepCompile, a compiler-based approach for systematically optimizing fully sharded training.